bill is 3 years younger than kylie their combined age is 49 how old is bill?

Answer by ikleyn(52788)

.

Bill + Kylie = 49


B    + (B+3) = 49

2B   +    3  = 49

2B           = 49-3 = 46

 B                  = 46/2 = 23.


ANSWER.  Bill is 23 years old.

Solved.